OpenStudy (anonymous):

I need some help on how to calculate base pairs There are 3,000,000,000 base pairs in mammalian genome. 10,000 genes in the mammalian genome A single gene average 10,000 base pairs in size 1. How many total base pairs are in all the mammalian genes? 2. What proportion (%) of the total genome does this represent?

OpenStudy (anonymous):

3.e+17 is the amount of total base pairs in all the mammalian genes

OpenStudy (aaronq):

1. gene bp = number of genes x average base pair size = = 10 000*10 000=100 000 000 =\(\sf 1.0*10^8\) 2. \(\sf \dfrac{gene~bp}{total~number~of~bp}*100\%=\dfrac{100~ 000~ 000}{ 3~000~000~000}*100\%=3.333\%\)